# Causal Classification of Pathological Misner-Type Spacetimes

### v4

@@ 4.1. Non-chronal region
In cylindrical coordinates $(\nu,r)$, the $\nu$–coordinate curves circle around the cylinder. Given the metric metric in Eddington-Finkelstein coordinates (as presented in Equation 4.4), besides $\theta,\phi$, we can also set $r$ to a constant value and consider the one-dimensional line element
@@ 5. Relation between pseudo-Schwarzschild and Misner spacetime
We aim to investigate Misner space and the 22–dimensional pseudo-Schwarzschild spacetime $(M_{Z},g_{Z})$, focusing on similarities such as global causal structure, curvature, and geodesics. In addition to these, the conformal structure of these spacetimes is notably rich. A conformal transformation can sometimes convert a curved space with a non-zero Riemann tensor into a flat space with a zero Riemann tensor. However, depending on the dimension of the spacetime, a less strict concept of causal relatedness may be required. In this context, we introduce the concept of isocausality [6], which refers to a situation where the causal structure of spacetime may not be preserved globally under certain transformations, but the notion of causal relationships within specific regions remains consistent.
A glance at the causal structure of the cylindrical pseudo-Schwarzschild space and the 22–dimensional Misner space might suggest that the two associated metrics, metric (4.4) and the 22–dimensional variant of metric (2.3), are related. This raises the natural question of whether an isometric mapping exists between these manifolds. However, it is known that any isometry of a pseudo-Riemannian manifold preserves its curvature. Since the Riemann curvature vanishes in Misner space, but does not vanish in the cylindrical pseudo-Schwarzschild space, no isometry exists between them.
@@
It is worth noting that in the special case $m=0$, the 22–dimensional pseudo-Schwarzschild spacetime is defined on $S^{1}\times\mathbb{R}$ and is flat. However, by setting $m=0$, the chronology horizon in the pseudo-Schwarzschild space disappears, and closed timelike curves (CTCs) no longer exist. This distinction makes these spacetimes causally different, as the chronology horizon is a significant global feature that separates the chronology-violating region from the region without CTCs. In what follows, we will restrict our investigation to the case where $m>0$.
@@ 5.1. General case
For radial trajectories in hyperbolically symmetric spacetimes, the geometry is effectively $1+1$ dimensional. Therefore, for simplicity and ease of calculation, we first focus on the 22–dimensional Misner space and the cylindrical pseudo-Schwarzschild space, as these lead to relatively straightforward conformal transformations. Afterward, we turn our attention to 44–dimensional spacetimes, although the results from the 22–dimensional case cannot be directly generalized. This gives rise to an alternative perspective on causal equivalence, as introduced by [6]. This approach allows for the consideration of a much larger class of causally related spacetimes compared to the classical concept of conformally related spacetimes.
@@ 5.1.1. Relation in the two-dimensional case
Since the metrics under study are 22–dimensional, we know that both metrics in question belong to the same conformal equivalence class $[g]$ of conformally flat metrics. [^10] A metric is said to be conformally flat if it can be conformally transformed into a metric with vanishing Riemann curvature. Hence, if the metrics $g_{P},g_{M}\in[g]$ are both locally conformal to a flat metric $g$, i.e. $(g_{M}\sim g)\,\wedge\,(g_{P}\sim g)$, then by transitivity and symmetry we have $g_{P}\sim g_{M}$. As a result, we can locally transform the curved pseudo-Schwarzschild space with a non-vanishing Riemann tensor into the flat Misner space with a zero Riemann tensor by applying a conformal transformation.
Consider the 22–dimensional flat Misner space defined on $S^{1}\times\mathbb{R}$ and equipped with the metric
@@
This version of Misner metric can be obtained from the 22–dimensional Minkowski metric (2.1) by the initial coordinate transformation $\tilde{t}=\eta\cosh(X_{1})$ and $\tilde{x}_{1}=\eta\sinh(X_{1})$. This coordinate change results in the Misner metric $ds^{2}=-d\eta^{2}+\eta^{2}(dX_{1})^{2}$, where $0<\eta<\infty$, $0\leq X_{1}\leq 2\pi$. Another coordinate change with $\varphi=X_{1}+\ln(\eta)$ and $T=\eta^{2}$ results in the Misner metric $dTd\varphi+Td\varphi{}^{2}$, where the domains are $\infty<T<\infty$ and $0\leq\varphi\leq 2\pi$.
@@
Conformal transformations between 22–dimensional Lorentzian manifolds send null hypersurfaces to null hypersurfaces and therefore null geodesics to null geodesics.
@@
This means that for the $1+1$ dimensional pseudo-Schwarzschild and Misner spacetimes, we can identify conformal transformations and demonstrate that timelike geodesics are of a similar type in both spacetimes. This establishes a notable relationship between them, with the 22–dimensional pseudo-Schwarzschild spacetime being viewed as a modification or generalization of Misner space. However, the 44–dimensional pseudo-Schwarzschild solution is not conformally flat, as indicated by the non-zero components of the conformally invariant Weyl tensor. In fact, only one independent spin component of the Weyl tensor does not vanish, specifically the Weyl scalar $C=\frac{m}{r^{3}}$. Therefore, in this context, a less strict definition of causal relatedness is required.
5.1.2. Relation in the 44–dimensional case
Since the Weyl curvature tensor vanishes identically in two dimensions, all 22–dimensional Lorentz metrics belong to the same conformal class of flat metrics. Therefore, it is relatively straightforward to find conformal transformations between any two representatives of this conformal class. In dimensions greater than three, the Weyl curvature tensor is generally non-zero, and a necessary condition for a spacetime to be conformally flat is that the Weyl tensor vanishes. As a result, establishing an explicit conformal relation between 44–dimensional spacetimes becomes significantly more challenging.
@@ 5.1.2. Relation in the 44–dimensional case
In earlier work [18] we conjectured that the 44–dimensional pseudo-Schwarzschild spacetime and Misner space are isocausal. In the present paper, we prove this result and extend it to include the pseudo-Reissner-Nordström spacetime. The proof, given in Proposition 7.1, establishes that all three models are pairwise isocausal on their universal covers and on suitable causally regular regions of the compactified spacetimes, and provides explicit criteria for when this equivalence descends globally.
@@ 6. Pseudo-Reissner-Nordström spacetime
Similar to the pseudo-Schwarzschild manifold, the 11–sphere $S^{1}$ (with $0\leq t<t+\alpha$) is topologically equivalent to an interval with its endpoints identified. The surface $H_{r}^{2}$ is 22–dimensional and corresponds to the upper sheet of a two-sheeted spacelike hyperboloid. The coordinate $\theta$ can take any positive value, while $\phi$ is periodic and ranges from $0\leqslant\phi<2\pi$. The real constants $m$ and $q$ represent mass and electric charge, respectively, and are assumed to be positive. This new metric can also be derived by performing a Wick rotation and a signature change on the Schwarzschild metric, followed by the substitution
@@
The Riemann curvature tensor for the 22–dimensional pseudo-Reissner-Nordström spacetime can be obtained using the same approach as in Section 4 for the pseudo-Schwarzschild case. The scalar curvature then follows immediately
@@
To obtain the curvature for the 44–dimensional case, we can derive the Riemann curvature tensor for the pseudo-Reissner-Nordström spacetime by applying a Wick rotation to the well-known Reissner-Nordström spacetime. This approach uses the established properties of the Reissner-Nordström solution, allowing us to modify the spacetime signature and adapt the curvature properties for the pseudo-Reissner-Nordström framework.
@@ 6.1.1. Non-chronal region $M_{2}$
Although the closed timelike curves (CTCs) in this case primarily arise from the periodicity of the $\nu$–coordinate, the tilting of light cones also contributes to the formation of CTCs.
@@ 6.2. Extremal pseudo-Reissner-Nordström spacetime
The extremal 22–dimensional pseudo-Reissner-Nordström spacetime $M$ with $m=q$ has only one horizon at $r=m$, which separates $M$ into two regions: $M_{3}=\{(\nu,r)\in M\mid 0<r<m\}$ and $M_{2}=\{(\nu,r)\in M\mid m<r\}$. As before, we time-orient $M$ by requiring $\partial_{r}$ to be future pointing. The metric degenerates to
@@ 6.3. Pseudo-Reissner-Nordström timelike geodesics
Due to the geodesic incompleteness, the embedded pseudo-Reissner-Nordström cylinder with $\theta=\textrm{const}$ and $\phi=\textrm{const}$, restricted to $M_{1}\cup M_{2}$, resembles the situation in the 22–dimensional Misner space and cylindrical pseudo-Schwarzschild spacetime.
@@ 7. Isocausality
with 22–dimensional Lorentzian base $(\mathcal{M}^{(Z)}_{i},g^{(Z)}_{i})$ admitting coordinates $(\nu,r)$ in which the metric takes the canonical Eddington-Finkelstein-type form
@@
Because every $1+1$ Lorentzian metric is locally conformally flat, any two bases $g^{(Z)}_{i}$ and $g^{(Z)}_{j}$ are locally related by a smooth, strictly positive conformal factor via a suitable change of null coordinates. By combining this with a careful control of the warp factor $r$, we obtain 44–dimensional causal maps that preserve the full light-cone structure (no restriction to base-only vectors). This observation motivates the main results of this section:
@@
Let $(\mathcal{M}_{\mathrm{M}},g_{\mathrm{M}})$ be Misner spacetime, $(\mathcal{M}_{\mathrm{PS}},g_{\mathrm{PS}})$ the pseudo-Schwarzschild spacetime, and $(\mathcal{M}_{\mathrm{PRN}},g_{\mathrm{PRN}})$ the pseudo-Reissner-Nordström spacetime, with the $1+1$ radial (base) metrics written in the canonical Eddington-Finkelstein-type form (7.1), and the full 44–metrics given by
@@
(B) Isocausality on causally regular regions. Let $U_{i}\subset\mathcal{M}_{i}$ denote open regions obtained by removing (a) neighborhoods of curvature singularities, and (b) neighborhoods of null hypersurfaces where $f_{i}=0$, so that $f_{i}$ is smooth and nowhere zero on $U_{i}$ and the $S^{1}$–orbits (if present) can be unwrapped. Then for every pair $(i,j)$ there exists a diffeomorphism $\Phi_{ij}:U_{i}\to U_{j}$ such that both $\Phi_{ij}$ and $\Phi_{ij}^{-1}$ preserve future-directed causal vectors; i.e. $(U_{i},g_{i})$ and $(U_{j},g_{j})$ are isocausal.
@@
Therefore a sufficient pointwise condition ensuring $\widetilde{\Phi}_{ij}$ is causal on the full 44–space is
@@
This is achieved for all sufficiently large $a$ (since the $a^{2}$-coefficients satisfy $2C_{1}^{2}<\frac{m_{j}}{M_{i}}v_{-}^{2}$ after possibly shrinking $I$; and any fixed constant term $2C_{0}^{2}$ is dominated for large $a$). Consequently, on any causally regular base interval $I$ bounded away from 0 and the zeros of $f_{i}$, we can choose $a>0$ (and solve (7.2)) so that (7.6)—hence (7.5)—holds. With this choice, $\widetilde{\Phi}_{ij}$ maps future-directed causal vectors for $g_{i}$ to future-directed causal vectors for $g_{j}$ in the full 44–dimensional warped product.
@@ 8. Conclusion
We have investigated a family of causality-violating spacetimes—specifically, the pseudo-Schwarzschild, pseudo-Reissner-Nordström, and Misner spacetimes—that share fundamental causal structures, including the presence of chronology horizons, Cauchy horizons, radial geodesics, and acausal regions. All these spacetimes can be expressed in a warped product form, with metrics decomposing into the cylindrical metric $g_{Z}$ and the hyperbolic metric $g_{H^{2}}$. This structure allows us to focus on their 22–dimensional cylindrical analogues. In Eddington-Finkelstein coordinates, these cylindrical versions can be brought to a canonical form $ds^{2}=g_{\nu\nu}d\nu^{2}+2g_{\nu r}d\nu dr$, from which the radial geodesic equations (Section 3) can be systematically derived. Renaming the coordinates as $\nu=\varphi$ and $r=T$, the connection to Misner spacetime becomes manifest.
@@
In each case, the base manifold $M_{Z}$ of the warped product structure is conformally related to a region of Minkowski spacetime and admits a Killing vector field whose orbits closely resemble those of the boost-generating vector field $X=x\,\partial_{t}+t\,\partial_{x}$ in two-dimensional Minkowski space. This Killing vector field generates a one-parameter group $G$ of isometries, some of whose orbits are timelike. A discrete subgroup $G_{0}\subset G$ acts properly and freely on $M_{Z}$, resulting in a Misner-like structure on the quotient $M_{Z}/G_{0}$, where the timelike orbits of $G$ project to closed timelike curves. [^11] As described in the classic book by Hawking and Ellis [11], Misner space is constructed by taking a quotient of half of Minkowski spacetime under a discrete subgroup of the one-parameter isometry group generated by the vector field $X=x\,\partial_{t}+t\,\partial_{x}$, which acts properly and freely in that region. A similar isometry group orbit structure appears in the (t,r)(t,r)–planes of both the Schwarzschild and Reissner–Nordström spacetimes, allowing for analogous quotient constructions in these cases as well. From a mathematical standpoint, this conformal structure renders the similarities in causal properties relatively unsurprising, as they directly reflect known features of Misner spacetime. Nevertheless, from a physics perspective, the results are noteworthy. The spacetimes examined differ markedly in their physical interpretation and geometric origin—ranging from a flat vacuum solution of the Einstein equations, to an asymptotically flat black hole solution of the vacuum Einstein equations, to a model that violates the weak energy condition and requires ‘exotic matter’ as its source—yet they exhibit analogous causal behavior. Understanding these similarities across physically distinct settings provides valuable insights into how causality can be broken or preserved under varying geometrical and physical conditions.
@@
In the context of ongoing research on spacetimes with closed timelike curves [1, 4, 9, 19], our findings offer a unified framework to view several causality-violating spacetimes through a common geometric and causal lens. The work reveals the intricate relationships between these examples and motivates a broader investigation into the interplay between geometry, causality, and spacetime topology. For instance, it would be compelling to investigate whether other spacetime models, such as the pseudo-Kerr spacetime [3]—a rotating generalization of the pseudo-Schwarzschild metric—also belong to this family of causality-violating, Misner-type spacetimes.
We view this article as a step toward a broader classification of causality-violating spacetimes. The warped-product ansatz and the explicit isocausality mappings presented here suggest that tools from conformal geometry and causal theory can provide a systematic framework for organising, comparing, and understanding causal structures in a wide range of gravitational models.

### v5

@@ 4.1. Non-chronal region
In cylindrical coordinates $(\nu,r)$, the $\nu$-coordinate curves circle around the cylinder. Given the metric metric in Eddington-Finkelstein coordinates (as presented in Equation 4.4), besides $\theta,\phi$, we can also set $r$ to a constant value and consider the one-dimensional line element
@@ 5. Relation between pseudo-Schwarzschild and Misner spacetime
We aim to investigate Misner space and the $2$-dimensional pseudo-Schwarzschild spacetime $(M_{Z},g_{Z})$, focusing on similarities such as global causal structure, curvature, and geodesics. In addition to these, the conformal structure of these spacetimes is notably rich. A conformal transformation can sometimes convert a curved space with a non-zero Riemann tensor into a flat space with a zero Riemann tensor. However, depending on the dimension of the spacetime, a less strict concept of causal relatedness may be required. In this context, we introduce the concept of isocausality [6], which refers to a situation where the causal structure of spacetime may not be preserved globally under certain transformations, but the notion of causal relationships within specific regions remains consistent.
A glance at the causal structure of the cylindrical pseudo-Schwarzschild space and the $2$-dimensional Misner space might suggest that the two associated metrics, metric (4.4) and the $2$-dimensional variant of metric (2.3), are related. This raises the natural question of whether an isometric mapping exists between these manifolds. However, it is known that any isometry of a pseudo-Riemannian manifold preserves its curvature. Since the Riemann curvature vanishes in Misner space, but does not vanish in the cylindrical pseudo-Schwarzschild space, no isometry exists between them.
@@
It is worth noting that in the special case $m=0$, the $2$-dimensional pseudo-Schwarzschild spacetime is defined on $S^{1}\times\mathbb{R}$ and is flat. However, by setting $m=0$, the chronology horizon in the pseudo-Schwarzschild space disappears, and closed timelike curves (CTCs) no longer exist. This distinction makes these spacetimes causally different, as the chronology horizon is a significant global feature that separates the chronology-violating region from the region without CTCs. In what follows, we will restrict our investigation to the case where $m>0$.
@@ 5.1. General case
For radial trajectories in hyperbolically symmetric spacetimes, the geometry is effectively $1+1$ dimensional. Therefore, for simplicity and ease of calculation, we first focus on the $2$-dimensional Misner space and the cylindrical pseudo-Schwarzschild space, as these lead to relatively straightforward conformal transformations. Afterward, we turn our attention to $4$-dimensional spacetimes, although the results from the $2$-dimensional case cannot be directly generalized. This gives rise to an alternative perspective on causal equivalence, as introduced by [6]. This approach allows for the consideration of a much larger class of causally related spacetimes compared to the classical concept of conformally related spacetimes.
@@ 5.1.1. Relation in the two-dimensional case
Since the metrics under study are $2$-dimensional, we know that both metrics in question belong to the same conformal equivalence class $[g]$ of conformally flat metrics. [^10] A metric is said to be conformally flat if it can be conformally transformed into a metric with vanishing Riemann curvature. Hence, if the metrics $g_{P},g_{M}\in[g]$ are both locally conformal to a flat metric $g$, i.e. $(g_{M}\sim g)\,\wedge\,(g_{P}\sim g)$, then by transitivity and symmetry we have $g_{P}\sim g_{M}$. As a result, we can locally transform the curved pseudo-Schwarzschild space with a non-vanishing Riemann tensor into the flat Misner space with a zero Riemann tensor by applying a conformal transformation.
Consider the $2$-dimensional flat Misner space defined on $S^{1}\times\mathbb{R}$ and equipped with the metric
@@
This version of Misner metric can be obtained from the $2$-dimensional Minkowski metric (2.1) by the initial coordinate transformation $\tilde{t}=\eta\cosh(X_{1})$ and $\tilde{x}_{1}=\eta\sinh(X_{1})$. This coordinate change results in the Misner metric $ds^{2}=-d\eta^{2}+\eta^{2}(dX_{1})^{2}$, where $0<\eta<\infty$, $0\leq X_{1}\leq 2\pi$. Another coordinate change with $\varphi=X_{1}+\ln(\eta)$ and $T=\eta^{2}$ results in the Misner metric $dTd\varphi+Td\varphi{}^{2}$, where the domains are $\infty<T<\infty$ and $0\leq\varphi\leq 2\pi$.
@@
Conformal transformations between $2$-dimensional Lorentzian manifolds send null hypersurfaces to null hypersurfaces and therefore null geodesics to null geodesics.
@@
This means that for the $1+1$ dimensional pseudo-Schwarzschild and Misner spacetimes, we can identify conformal transformations and demonstrate that timelike geodesics are of a similar type in both spacetimes. This establishes a notable relationship between them, with the $2$-dimensional pseudo-Schwarzschild spacetime being viewed as a modification or generalization of Misner space. However, the $4$-dimensional pseudo-Schwarzschild solution is not conformally flat, as indicated by the non-zero components of the conformally invariant Weyl tensor. In fact, only one independent spin component of the Weyl tensor does not vanish, specifically the Weyl scalar $C=\frac{m}{r^{3}}$. Therefore, in this context, a less strict definition of causal relatedness is required.
5.1.2. Relation in the $4$-dimensional case
Since the Weyl curvature tensor vanishes identically in two dimensions, all $2$-dimensional Lorentz metrics belong to the same conformal class of flat metrics. Therefore, it is relatively straightforward to find conformal transformations between any two representatives of this conformal class. In dimensions greater than three, the Weyl curvature tensor is generally non-zero, and a necessary condition for a spacetime to be conformally flat is that the Weyl tensor vanishes. As a result, establishing an explicit conformal relation between $4$-dimensional spacetimes becomes significantly more challenging.
@@ 5.1.2. Relation in the $4$-dimensional case
In earlier work [18] we conjectured that the $4$-dimensional pseudo-Schwarzschild spacetime and Misner space are isocausal. In the present paper, we prove this result and extend it to include the pseudo-Reissner-Nordström spacetime. The proof, given in Proposition 7.1, establishes that all three models are pairwise isocausal on their universal covers and on suitable causally regular regions of the compactified spacetimes, and provides explicit criteria for when this equivalence descends globally.
@@ 6. Pseudo-Reissner-Nordström spacetime
Similar to the pseudo-Schwarzschild manifold, the $1$-sphere $S^{1}$ (with $0\leq t<t+\alpha$) is topologically equivalent to an interval with its endpoints identified. The surface $H_{r}^{2}$ is $2$-dimensional and corresponds to the upper sheet of a two-sheeted spacelike hyperboloid. The coordinate $\theta$ can take any positive value, while $\phi$ is periodic and ranges from $0\leqslant\phi<2\pi$. The real constants $m$ and $q$ represent mass and electric charge, respectively, and are assumed to be positive. This new metric can also be derived by performing a Wick rotation and a signature change on the Schwarzschild metric, followed by the substitution
@@
The Riemann curvature tensor for the $2$-dimensional pseudo-Reissner-Nordström spacetime can be obtained using the same approach as in Section 4 for the pseudo-Schwarzschild case. The scalar curvature then follows immediately
@@
To obtain the curvature for the $4$-dimensional case, we can derive the Riemann curvature tensor for the pseudo-Reissner-Nordström spacetime by applying a Wick rotation to the well-known Reissner-Nordström spacetime. This approach uses the established properties of the Reissner-Nordström solution, allowing us to modify the spacetime signature and adapt the curvature properties for the pseudo-Reissner-Nordström framework.
@@ 6.1.1. Non-chronal region $M_{2}$
Although the closed timelike curves (CTCs) in this case primarily arise from the periodicity of the $\nu$-coordinate, the tilting of light cones also contributes to the formation of CTCs.
@@ 6.2. Extremal pseudo-Reissner-Nordström spacetime
The extremal $2$-dimensional pseudo-Reissner-Nordström spacetime $M$ with $m=q$ has only one horizon at $r=m$, which separates $M$ into two regions: $M_{3}=\{(\nu,r)\in M\mid 0<r<m\}$ and $M_{2}=\{(\nu,r)\in M\mid m<r\}$. As before, we time-orient $M$ by requiring $\partial_{r}$ to be future pointing. The metric degenerates to
@@ 6.3. Pseudo-Reissner-Nordström timelike geodesics
Due to the geodesic incompleteness, the embedded pseudo-Reissner-Nordström cylinder with $\theta=\textrm{const}$ and $\phi=\textrm{const}$, restricted to $M_{1}\cup M_{2}$, resembles the situation in the $2$-dimensional Misner space and cylindrical pseudo-Schwarzschild spacetime.
@@ 7. Isocausality
with $2$-dimensional Lorentzian base $(\mathcal{M}^{(Z)}_{i},g^{(Z)}_{i})$ admitting coordinates $(\nu,r)$ in which the metric takes the canonical Eddington-Finkelstein-type form
@@
Because every $1+1$ Lorentzian metric is locally conformally flat, any two bases $g^{(Z)}_{i}$ and $g^{(Z)}_{j}$ are locally related by a smooth, strictly positive conformal factor via a suitable change of null coordinates. By combining this with a careful control of the warp factor $r$, we obtain $4$-dimensional causal maps that preserve the full light-cone structure (no restriction to base-only vectors). This observation motivates the main results of this section:
@@
Let $(\mathcal{M}_{\mathrm{M}},g_{\mathrm{M}})$ be Misner spacetime, $(\mathcal{M}_{\mathrm{PS}},g_{\mathrm{PS}})$ the pseudo-Schwarzschild spacetime, and $(\mathcal{M}_{\mathrm{PRN}},g_{\mathrm{PRN}})$ the pseudo-Reissner-Nordström spacetime, with the $1+1$ radial (base) metrics written in the canonical Eddington-Finkelstein-type form (7.1), and the full $4$-metrics given by
@@
(B) Isocausality on causally regular regions. Let $U_{i}\subset\mathcal{M}_{i}$ denote open regions obtained by removing (a) neighborhoods of curvature singularities, and (b) neighborhoods of null hypersurfaces where $f_{i}=0$, so that $f_{i}$ is smooth and nowhere zero on $U_{i}$ and the $S^{1}$-orbits (if present) can be unwrapped. Then for every pair $(i,j)$ there exists a diffeomorphism $\Phi_{ij}:U_{i}\to U_{j}$ such that both $\Phi_{ij}$ and $\Phi_{ij}^{-1}$ preserve future-directed causal vectors; i.e. $(U_{i},g_{i})$ and $(U_{j},g_{j})$ are isocausal.
@@
Therefore a sufficient pointwise condition ensuring $\widetilde{\Phi}_{ij}$ is causal on the full $4$-space is
@@
This is achieved for all sufficiently large $a$ (since the $a^{2}$-coefficients satisfy $2C_{1}^{2}<\frac{m_{j}}{M_{i}}v_{-}^{2}$ after possibly shrinking $I$; and any fixed constant term $2C_{0}^{2}$ is dominated for large $a$). Consequently, on any causally regular base interval $I$ bounded away from 0 and the zeros of $f_{i}$, we can choose $a>0$ (and solve (7.2)) so that (7.6)—hence (7.5)—holds. With this choice, $\widetilde{\Phi}_{ij}$ maps future-directed causal vectors for $g_{i}$ to future-directed causal vectors for $g_{j}$ in the full $4$-dimensional warped product.
@@ 8. Conclusion
We have investigated a family of causality-violating spacetimes—specifically, the pseudo-Schwarzschild, pseudo-Reissner-Nordström, and Misner spacetimes—that share fundamental causal structures, including the presence of chronology horizons, Cauchy horizons, radial geodesics, and acausal regions. All these spacetimes can be expressed in a warped product form, with metrics decomposing into the cylindrical metric $g_{Z}$ and the hyperbolic metric $g_{H^{2}}$. This structure allows us to focus on their $2$-dimensional cylindrical analogues. In Eddington-Finkelstein coordinates, these cylindrical versions can be brought to a canonical form $ds^{2}=g_{\nu\nu}d\nu^{2}+2g_{\nu r}d\nu dr$, from which the radial geodesic equations (Section 3) can be systematically derived. Renaming the coordinates as $\nu=\varphi$ and $r=T$, the connection to Misner spacetime becomes manifest.
@@
In each case, the base manifold $M_{Z}$ of the warped product structure is conformally related to a region of Minkowski spacetime and admits a Killing vector field whose orbits closely resemble those of the boost-generating vector field $X=x\,\partial_{t}+t\,\partial_{x}$ in two-dimensional Minkowski space. This Killing vector field generates a one-parameter group $G$ of isometries, some of whose orbits are timelike. A discrete subgroup $G_{0}\subset G$ acts properly and freely on $M_{Z}$, resulting in a Misner-like structure on the quotient $M_{Z}/G_{0}$, where the timelike orbits of $G$ project to closed timelike curves. [^11] As described in the classic book by Hawking and Ellis [11], Misner space is constructed by taking a quotient of half of Minkowski spacetime under a discrete subgroup of the one-parameter isometry group generated by the vector field $X=x\,\partial_{t}+t\,\partial_{x}$, which acts properly and freely in that region. A similar isometry group orbit structure appears in the $(t,r)$-planes of both the Schwarzschild and Reissner-Nordström spacetimes, allowing for analogous quotient constructions in these cases as well. From a mathematical standpoint, this conformal structure renders the similarities in causal properties relatively unsurprising, as they directly reflect known features of Misner spacetime. Nevertheless, from a physics perspective, the results are noteworthy. The spacetimes examined differ markedly in their physical interpretation and geometric origin—ranging from a flat vacuum solution of the Einstein equations, to an asymptotically flat black hole solution of the vacuum Einstein equations, to a model that violates the weak energy condition and requires “exotic matter” as its source—yet they exhibit analogous causal behavior. Understanding these similarities across physically distinct settings provides valuable insights into how causality can be broken or preserved under varying geometrical and physical conditions.
While certain models considered here require violations of classical energy conditions, their study remains valuable as a probe of the causal structure permitted by general relativity. Specifically, the observation that vacuum and non-vacuum models can exhibit identical causal behavior suggests that these pathological features are fundamentally geometric rather than artifacts of specific matter content.
@@
In the context of ongoing research on spacetimes with closed timelike curves [1, 4, 9, 19], our findings offer a unified framework to view several causality-violating spacetimes through a common geometric and causal lens. The work reveals the intricate relationships between these examples and motivates a broader investigation into the interplay between geometry, causality, and spacetime topology. For instance, it would be compelling to investigate whether other spacetime models, such as the pseudo-Kerr spacetime [3]—a rotating generalization of the pseudo-Schwarzschild metric—also belong to this family of causality-violating, Misner-type spacetimes. We view this article as a step toward a broader classification of causality-violating spacetimes. The warped-product ansatz and the explicit isocausality mappings presented here suggest that tools from conformal geometry and causal theory can provide a systematic framework for organising, comparing, and understanding causal structures in a wide range of gravitational models.